Circular No. 6848

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SAGITTARII 1998 M. Della Valle, Astronomy Department, Padua University; and A.

Pizzella and M. Bernardi, European Southern Observatory, report: “Preliminary analysis of a spectrum of N Sgr 1998 (IAUC 6846, 6847), obtained at La Silla with the 1.54-m Danish telescope (+ Danish Faint Object Spectrograph and Camera; range 370–760 nm, resolution ∼ 1 nm) on Mar. 24.3 UT, confirms this object to be a galactic nova caught during an early stage. The continuum appears dominated by emission lines of the Balmer series, flanked by weak P-Cyg profiles, Fe II (multiplets 42, 48, 49, 74, 224), and N II and Na I lines. The massive presence of iron and the relatively slow faint expansion velocities inferred from the FWHMs of Hα and Hβ (1600 and 1800 km/s, respectively) suggest that this nova may belong to the Fe II class (see Williams 1992, A.J. 104, 725).”

Visual magnitude estimates: Mar. 23.687 UT, 7.4 (A. F. Jones, Nelson, New Zealand); 24.12, 8.3 (L. Kiss, Szeged, Hungary); 24.139, 8.1 (K. Hornoch, Lelekovice, Czech Republic); 24.183, 8.5 (P. Schmeer, Bischmisheim, Germany); 24.21, 7.6 (J. M. Trigo, Benicassim, Spain).

GRB 970508 A. J. Castro-Tirado and J. Gorosabel, Laboratorio de Astrofisica Es-

pacial y Fisica Fundamental, Madrid; T. Galama and P. Groot, University of Amsterdam; J. van Paradijs, University of Amsterdam and University of Alabama at Hunstville; and C. Kouveliotou, Universities Space Research Association, report: “R-band images have been obtained by J. H. Telting, B. Garcia-Lorenzo, and L. Jones (Isaac Newton Group Telescopes) with the 4.2-m William Herschel Telescope on La Palma on Mar. 20/21. At the same position of the optical counterpart to GRB 970508 reported by Bond (IAUC 6654), the co-added image (13700 s) reveals an object 5.4 mag fainter: R = 25.20 ± 0.25 (based on the secondary standards given by

Sokolov et al. 1998, at http://xxx.lanl.gov/abs/astro-ph/?9802341).

This is ∼ 1.0 mag brighter than expected from the extrapolation of the power-law decline (R = 26.18 ± 0.09), suggesting that the optical light is dominated by the host galaxy. Assuming its location at z = 0.8 (Metzger et al. 1997, Nature 387, 879), this corresponds to an absolute blue magnitude of −17.9 ± 0.5, placing the object below the knee of the galaxy luminosity function.”

1998 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green